OpenStudy (anonymous):

Look at the pattern shown below. What number is missing from the pattern? 20, 22, 26, 34, __,82

OpenStudy (anonymous):

the difference is doubling

OpenStudy (anonymous):

2,4,8,16 20+2=22 22+4=26 26+8=34 34+16=50 50+32=82

OpenStudy (anonymous):

ok wow thx